OpenStudy (anonymous):

g(x) = x2 + 3, what is the change made to g(x) if the parent function is f(x) = x2? Yeah totally stumped here.

OpenStudy (anonymous):

this isn't as hard as you think...

OpenStudy (anonymous):

Im over thinking it.. i know :/

OpenStudy (anonymous):

if I have a function f(x) = x^2, and you want to make me have a function that is "x^2 + 3" what do you need to do to my function?

OpenStudy (anonymous):

im looking haha

OpenStudy (anonymous):

they are practically the same, right? but you wish mine had a "plus 3" on the end... so that's it... g(x) = f(x) + 3 because f(x) is just x^2 So g(x) = x^2 + 3

OpenStudy (anonymous):

Try graphing the functions and you will probably see the change. Or try making a t-table with both functions and see if you can see what change is the result.

OpenStudy (anonymous):

so would it be a vertical translation of 3 units up?

OpenStudy (anonymous):

unless I'm reading "parent function" wrong, I think that's all they're doing here... it's a question like your last one that shows you the impact of changing a function somehow, in this case, by adding 3

OpenStudy (anonymous):

Yes, nice!

OpenStudy (anonymous):